Refer to the National Income Accounting Matrix Table(in Moodle)and use it to fill in the values of the variables in the matrixcalculatedfrom the Keynesian algebraic model given below. a.Work throughthe problem belowand show working for all variables(1-10) in addition to totals of columns and rows in the matrix.(10 marks)C = 300 + 0.8YD (Private consumption)YD = Y + TR –T (Private Disposable income)T = 80 + 0.1Y (Total taxes)G = 300, TR = 100 (Gov. expenditure and transfer payments respectively)I = 450 (Private investment)X = 200 (Total exports)IM = 120 + 0.2Y (Total imports)Y = C + I + G + X –M (Goods Market Equilibrium condition)Policy variables: Fiscal policy (G, t and TR), Monetary Policy: (N/A)
1
Expert's answer
2019-05-14T09:00:48-0400

Y = C + I + G + X – M = 300 + 0.8×(Y + 100 - 80 - 0.1Y) + 450 + 300 + 200 - 120 - 0.2Y,

Y - 0.72Y + 0.2Y = 300 + 16 + 450 + 300 + 200 - 120,

0.48Y = 1146,

Y = 2387.5.